Circular No. 8341

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004bs AND 2004bt Further to IAUC 8336, M. Armstrong reports his discovery of an appar-

ent supernova (mag 17.0) on unfiltered CCD images taken on May 16.914 and 17.884 UT. SN 2004bs is located at α = 10^{h}39^{m}39^{s}.30, δ = +25^{o}19′19′′.9 (equinox 2000.0), which is ≈ 3′′.4 east and 1′′.7 south of the center of NGC 3323. Nothing was visible at this position on his images taken on 2000 Dec. 22 (limiting mag ∼ 19.5), 2004 Apr. 13, 19, and 25 (limiting mag ∼ 19) or on Palomar Sky survey red and blue plates.

Further to IAUC 8311, T. Boles reports the discovery of an apparent supernova (mag 18.2) on unfiltered CCD images taken on May 17.023 and 18.001 UT. The new object is located at α = 14^{h}19^{m}53^{s}.88, δ = +51^{o}53′49′′.6, which is ≈ 6′′.3 west and 8′′.1 north of the center of UGC 9178. SN 2004bt is not present on Boles’ images from Mar. 22 and Feb. 20 (limiting mag 19.5), and it is not present on Digitized Sky Survey red (1994) or blue (1995) plates.

COMETS C/2003 Q10, T9–T11, U5, U6, W2, Y1 (SOHO) Further to IAUC 8340, K. Battams reports measurements for addi-

tional near-sun comets found on SOHO website images by M. Boschat (C/2003 Q10, T11), X.-m. Zhou (C/2003 T9, W2), J. Sachs (C/2003 T10), X. Leprette (C/2003 U5), and M. Meyer (C/2003 Y1). C/2003 Q10 was observed only in C3 images, while C/2003 T9, T11, U5, and W2 were observed only in C2 images; the remaining comets were observed in both coronagraphs. The reduced astrometry by B. G. Marsden and R. Kisala, and orbital elements by Marsden, appear on the MPECs cited below. C/2003 W2 and C/2003 Y1 belong to the Meyer group; the remaining objects are Kreutz sungrazers.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 Q10 Aug. 25.113  9^{h}57.^{m}0  +10^{o}29′  2004-K16
C/2003 T9  Oct. 9.746  12 51.1  − 6 48  2004-K16
C/2003 T10 Oct. 13.404  12 49.8  − 9 16  2004-K16
C/2003 T11 Oct. 15.479  13 12.5  − 9 12  2004-K16
C/2003 U5 Oct. 16.996  13 18.0  − 9 46  2004-K16
C/2003 U6 Oct. 20.638  13 20.9  −12 04  2004-K16
C/2003 W2 Nov. 29.379  16 19.3  −23 07  2004-J21
C/2003 Y1 Dec. 20.379  17 52.3  −22 21  2004-J21
2004 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green